# From Lengthy to Lucid: A Systematic Literature Review on NLP Techniques for Taming Long Sentences

### v1

@@ 4.3. Evaluation Metrics
(3) $BP=\left\{\begin{array}[]{cl}1&\textrm{if }c>r\\ \exp^{(1-r/c)}&\textrm{if }c\leq r\end{array}\right.$
@@
(4) $p_{n}=\frac{\sum\limits_{C\in\{Candidates\}}\sum\limits_{\textrm{n-gram}\in C}Count_{clip}{(\textrm{n-gram})}}{\sum\limits_{C^{\prime}\in\{Candidates\}}\sum\limits_{\textrm{n-gram'}\in C^{\prime}}Count(\textrm{n-gram'})}$
, where $Candidates$ represent the set of generated sentences, $count$ indicates the count of n-grams in the candidate sentence, and $Count_{clip}=min(count,max\_ref\_count)$ with $\max\_ref\_count$ used to limit count to the maximum count of the n-gram in the reference candidate.
While BLEU can serve as an indicator of the quality of the generated text according to the ground truth, it is not capable of taking into account the semantics of a sentence as other metrics, i.e., BERTScore (Zhang et al., 2020a). BERTScore, which is based on the Transformer architecture, is widely used to measure the quality of generated text in various natural language generation tasks, such as machine translation, summarization, etc. Unlike BLEU, which relies on simple n-gram overlaps, BERTScore exploits BERT’s contextual embeddings. To compute the similarity between a ground truth and a predicted set of sentences, the cosine similarities of each token embedding of the input sentences are employed. More specifically, BERTscore calculates the recall and precision for each token embedding in the generated and target sentences. BERTScore then aggregates the F1 scores to provide an overall similarity score that is used to measure how well the generated text reflects the ground truth as follows:
(5) $BERT_{score}=\frac{2\cdot P_{BERT}\cdot R_{BERT}}{P_{BERT}+R_{BERT}},$
@@
Several works also report the ratio of simple sentences per complex sentence (#$\#S/C$), as well as the ratio of tokens per simple sentence (#$\#T/S$). The #$\#S/C$ ratio can give a rough estimation of the complexity of the given pair of sentences. A higher #$\#S/C$ ratio indicates a higher proportion of simple sentences. Similarly, the #$\#T/S$ score represents the ratio of the average number of tokens per simple sentence. A low #$\#T/S$ ratio indicates that fewer tokens are included in simple sentences while sentences with higher scores tend to contain more tokens. These ratios provide valuable insights into both the complexity and the structure of the generated sentences. Typically, these metrics are computed in relation to the ground truth sentences within a dataset.
4.4. Comparative Evaluation
@@
Table 7. Results of splitting methods on different datasets. BLEU, BertScore, #$\#S/C$ and #$\#T/S$ are reported. Results are obtained from the corresponding papers. Dashes (-) indicate that the score is not provided.
@@
(Aharoni and Goldberg, 2018) 25.47 - 2.3 11.8
(Botha et al., 2018) 30.5 - 2.0 8.8
(Zhang et al., 2020b) (LSTM) 62.6 - - -
(Zhang et al., 2020b) (rules) 65.9 - - -
(Fan et al., 2020) 71.2 - 2.5 10.3
Wiki-BM (Zhang et al., 2020b) (rules) 77.2 - - -
(Zhang et al., 2020b) (LSTM) 87.0 - - -
BiSECT (Kim et al., 2021) 45.8 85.6 - -
(Aharoni and Goldberg, 2018) 46.3 85.3 - -
Regarding the WebSplit dataset, we notice that the use of simple rules (Zhang et al., 2020b) outperforms all the RNN-based models with a BLEU score of 65.9% compared to 62.6% of the best performing LSTM model. This finding can be attributed to the simple and repetitive patterns of WebSplit as highlighted in (Zhang et al., 2020b). On the other hand, the Transformer memory-based model of (Fan et al., 2020) achieves the best results in terms of BLEU score and the closest #$\#S/C$ and #$\#T/C$ ratios to the reference ones.
Regarding the Wiki-BM test set, the best performance is obtained using an LSTM with attention and a copying mechanism (Zhang et al., 2020b). Note that this model is trained on the WikiSplit dataset. Unlike WebSplit, the rule-based method does not perform as well on Wiki-BM, since it is a more challenging dataset with more complex examples.
For the BiSECT corpus, both the RNN-based model (Aharoni and Goldberg, 2018) and the Transformer-based model (Kim et al., 2021) achieve comparable results ($\sim 46\%$ in terms of BLEU and $\sim 85\%$ in terms of BERTScore).
In conclusion, the overall comparative evaluation reveals an interesting trend regarding the rule-based methods when dealing with simpler datasets such as WebSplit. On the other hand, these methods struggle to perform when more complex datasets such as Wiki-BM are introduced. Transformer and LSTM models consistently achieve the best performance across the different splitting datasets.
5. Challenges and Limitations
@@
5.1. Sentence Compression vs Sentence Splitting
@@
5.2. Evaluation Metrics
Despite the significant progress on both sentence compression and sentence splitting methods, the field lacks an established automated evaluation metric that is capable of measuring the meaning preservation between the generated compression and the original long sentence. While several evaluation metrics exist to evaluate the quality of the resolved sentences, such as ROUGE, BLUE or grammatical F1 score, these are focused only on the overlap of tokens as well as grammatical or syntactic consistency. The evaluation of meaning preservation is typically assessed only by human annotators, which significantly limits its scalability and introduces additional effort and costs. In fact, there is only one work (Beauchemin et al., 2023) for evaluating meaning preservation within the field of sentence simplification, but currently there are no established metrics for sentence compression and sentence splitting for assessing this critical aspect. The importance of a meaning preservation metric becomes evident when we consider the primary objective of resolving a long sentence. By retaining the original meaning of a sentence while shortening or splitting it, we ensure that the resulting text will remain logically coherent and will deliver accurately the intended message. This is especially important in contexts where preserving the meaning is crucial, such as news articles, social media posts or legal documents. Such a metric can significantly advance research and benefit practical applications in sentence compression and sentence splitting.
5.3. Interpretability
@@ 5.4. Complexity
6. Conclusions
This paper provided a systematic review of sentence compression and splitting methods for resolving long sentences. We covered a wide spectrum of techniques, starting from traditional approaches, such as rule-based and statistical methods, and progressing towards more advanced solutions, such as RNNs and Transformers models. We systematically categorized these methods into a comprehensive taxonomy and we provided a comparative analysis between existing models on widely used datasets for both tasks. In addition, we discussed the challenges and limitations associated with evaluating the effectiveness of these techniques. Supervised techniques, especially extractive ones, are the most commonly employed for sentence compression. With the advent of deep learning, there has been a significant increase in the development of abstractive sentence compression methods. Similarly, for sentence splitting, the landscape is dominated by supervised methods, with Transformers and RNN-based methods being the most effective ones. However, in simple scenarios, rule-based methods have also demonstrated their ability to recognize straightforward patterns. We observed a significant gap in weakly supervised and self-supervised techniques for sentence compression and unsupervised methods for sentence splitting. These unexplored areas provide the opportunity for further research on methods that do not rely on large annotated datasets, offering potential solutions to domains with limited data.

### v2

@@ 4.3 Evaluation Metrics
$BP=\left\{\begin{array}[]{cl}1&\textrm{if }c>r\\ \exp^{(1-r/c)}&\textrm{if }c\leq r\end{array}\right$. (3)
@@
$p_{n}=\frac{\sum\limits_{C\in\{Candidates\}}\sum\limits_{\textrm{n-gram}\in C}Count_{clip}{(\textrm{n-gram})}}{\sum\limits_{C^{\prime}\in\{Candidates\}}\sum\limits_{\textrm{n-gram'}\in C^{\prime}}Count(\textrm{n-gram'})}$ (4)
, where $Candidates$ represent the set of generated sentences, countcount indicates the count of n-grams in the candidate sentence, and $Count_{clip}=min(count,max\_ref\_count)$ with $\max\_ref\_count$ used to limit count to the maximum count of the n-gram in the reference candidate.
While BLEU can serve as an indicator of the quality of the generated text according to the ground truth, it is not capable of taking into account the semantics of a sentence as other metrics, i.e., BERTScore Zhang2020BertScore . BERTScore, which is based on the Transformer architecture, is widely used to measure the quality of generated text in various natural language generation tasks, such as machine translation, summarization, etc. Unlike BLEU, which relies on simple n-gram overlaps, BERTScore exploits BERT’s contextual embeddings. To compute the similarity between a ground truth and a predicted set of sentences, the cosine similarities of each token embedding of the input sentences are employed. More specifically, BERTscore calculates the recall and precision for each token embedding in the generated and target sentences. BERTScore then aggregates the F1 scores to provide an overall similarity score that is used to measure how well the generated text reflects the ground truth as follows:
$BERT_{score}=\frac{2\cdot P_{BERT}\cdot R_{BERT}}{P_{BERT}+R_{BERT}},$ (5)
@@
Several works also report the ratio of simple sentences per complex sentence (#S/C\#S/C), as well as the ratio of tokens per simple sentence (#T/S\#T/S). The #S/C\#S/C ratio can give a rough estimation of the complexity of the given pair of sentences. A higher #S/C\#S/C ratio indicates a higher proportion of simple sentences. Similarly, the #T/S\#T/S score represents the ratio of the average number of tokens per simple sentence. A low #T/S\#T/S ratio indicates that fewer tokens are included in simple sentences while sentences with higher scores tend to contain more tokens. These ratios provide valuable insights into both the complexity and the structure of the generated sentences. Typically, these metrics are computed in relation to the ground truth sentences within a dataset.
4.4 Comparative Evaluation
@@
Table 7: Results of splitting methods on different datasets. BLEU, BertScore, #S/C\#S/C and #T/S\#T/S are reported. Results are obtained from the corresponding papers. Dashes (-) indicate that the score is not provided.
@@
aharoni-goldberg-2018-split 25.47 - 2.3 11.8
botha-etal-2018-learning 30.5 - 2.0 8.8
Zhang2020 (LSTM) 62.6 - - -
Zhang2020 (rules) 65.9 - - -
fan2020memory 71.2 - 2.5 10.3
Wiki-BM ponce-etal-2024-split 74.40 - - -
Zhang2020 (rules) 77.2 - - -
Zhang2020 (LSTM) 87.0 - - -
BiSECT ponce-etal-2024-split 42.4 - - -
kim-etal-2021-bisect 45.8 85.6 - -
aharoni-goldberg-2018-split 46.3 85.3 - -
Regarding the WebSplit dataset, we notice that the use of simple rules Zhang2020 outperforms all the RNN-based models with a BLEU score of 65.9% compared to 62.6% of the best performing LSTM model. This finding can be attributed to the simple and repetitive patterns of WebSplit as highlighted in Zhang2020 . On the other hand, the Transformer memory-based model of fan2020memory achieves the best results in terms of BLEU score and the closest #S/C\#S/C and #T/C\#T/C ratios to the reference ones.
Regarding the Wiki-BM test set, the best performance is obtained using an LSTM with attention and a copying mechanism Zhang2020 . Note that this model is trained on the WikiSplit dataset. Unlike WebSplit, the rule-based method does not perform as well on Wiki-BM, since it is a more challenging dataset with more complex examples.
For the BiSECT corpus, both the RNN-based model aharoni-goldberg-2018-split and the Transformer-based model kim-etal-2021-bisect achieve comparable results ($\sim 46\%$ in terms of BLEU and $\sim 85\%$ in terms of BERTScore).
In conclusion, the overall comparative evaluation reveals an interesting trend regarding the rule-based methods when dealing with simpler datasets such as WebSplit. On the other hand, these methods struggle to perform when more complex datasets such as Wiki-BM are introduced. Transformer and LSTM models consistently achieve the best performance across the different splitting datasets. In addition, LLMs did not outperform traditional neural models in either the Wiki-BM or BiSECT test sets. Interestingly, this limitation suggests a promising research direction for further enhancing LLMs to better handle these tasks.
5 Challenges and Limitations
@@
5.1 Sentence Compression vs Sentence Splitting
@@
5.2 Evaluation Metrics
Despite the significant progress on both sentence compression and sentence splitting methods, the field lacks an established automated evaluation metric that is capable of measuring the meaning preservation between the generated compression and the original long sentence. While several evaluation metrics exist to evaluate the quality of the resolved sentences, such as ROUGE, BLUE or grammatical F1 score, these are focused only on the overlap of tokens as well as grammatical or syntactic consistency. The evaluation of meaning preservation is typically assessed only by human annotators, which significantly limits its scalability and introduces additional effort and costs. In fact, there is only one work Beauchemin2023Meaning for evaluating meaning preservation within the field of sentence simplification, but currently there are no established metrics for sentence compression and sentence splitting for assessing this critical aspect. The importance of a meaning preservation metric becomes evident when we consider the primary objective of resolving a long sentence. By retaining the original meaning of a sentence while shortening or splitting it, we ensure that the resulting text will remain logically coherent and will deliver accurately the intended message. This is especially important in contexts where preserving the meaning is crucial, such as news articles, social media posts or legal documents. Such a metric can significantly advance research and benefit practical applications in sentence compression and sentence splitting.
5.3 Interpretability
@@ 5.4 Complexity
5.5 Large Language Models in Sentence Compression and Splitting
Despite the impressive performance of LLMs in various NLP tasks, their application in sentence compression and splitting has been limited. LLMs have the potential to significantly improve these tasks by providing deeper contextual understanding and enhancing adaptability to new domains and languages. There is strong reason to expect that LLMs could excel in sentence compression and splitting, particularly given their demonstrated success in abstractive text summarization liu2023revisiting ; zhang2024benchmarking . Since sentence compression can be viewed as an extremely fine-grained form of summarization at the sentence level, and sentence splitting as a controlled form of content expansion and reorganization, the mechanisms used in summarization are directly applicable to both tasks.
Given this alignment, LLMs could substantially improve sentence compression and splitting by enabling more accurate decisions about which sentence elements are essential and how clauses should be reorganized. Furthermore, because LLMs generalize well across domains, they can support compression tasks in low-resource or specialized areas by generating high-quality synthetic training data or by acting as teacher models in distillation pipelines Raffel2020 . This would allow lightweight models to retain the benefits of LLM-level understanding while remaining feasible for real-time or resource-constrained deployment.
However, given the computational demands and environmental impacts of these models, applying LLMs in this area requires a balance between their capabilities and resource limitations. For example, LLMs could be used strategically to generate training samples for low-resource domains or languages, while lighter models could be used for fine-tuning in these tasks. Additionally, multimodal sentence compression, which integrates information from both text and images, may particularly benefit from the contextual capabilities of LLMs, though achieving this efficiently remains an ongoing research challenge.
6 Conclusions
This paper provided a systematic review of sentence compression and splitting methods for resolving long sentences. We covered a wide spectrum of techniques, starting from traditional approaches, such as rule-based and statistical methods, and progressing towards more advanced solutions, such as RNNs, Transformers models and LLMs. We systematically categorized these methods into a comprehensive taxonomy and we provided a comparative analysis between existing models on widely used datasets for both tasks. In addition, we discussed the challenges and limitations associated with evaluating the effectiveness of these techniques. Supervised techniques, especially extractive ones, are the most commonly employed for sentence compression. With the advent of deep learning, there has been a significant increase in the development of abstractive sentence compression methods. Similarly, for sentence splitting, the landscape is dominated by supervised methods, with Transformers and RNN-based methods being the most effective ones. However, in simple scenarios, rule-based methods have also demonstrated their ability to recognize straightforward patterns. We observed a significant gap in weakly supervised and self-supervised techniques for sentence compression and unsupervised methods for sentence splitting. We also noted that, despite their potential, LLMs have not yet been extensively studied for these tasks. These unexplored areas provide the opportunity for further research on methods that do not rely on large annotated datasets, offering potential solutions to domains with limited data.